OpenStudy (anonymous):

The value of acceleration due to gravity (g) on Pluto is about 0.61 meters/second2. How much will an object that weighs 250 newtons on Earth weigh on Pluto? Note that the value of acceleration due to gravity on Earth is 9.8 meters/second2.

OpenStudy (ganpat):

on Earth = Weight = Mass * g earth So Mass of object = 250/ 9.8 = 25.51 Kg.. On Pluto ' Weight = mass * g pluto = 25.52 * 0.61 = 15.3 N...

OpenStudy (anonymous):

\[F=ma\] 1 (kg)=9.8 (N) so 250 (N)=25.5 (kg) then, in Pluto \[F=ma=25.5\times0.61=15.56 (N)\]